# Do we still need Human Annotators? Prompting Large Language Models for Aspect Sentiment Quad Prediction

### v2

@@ E.2 ASQP
Aspect Term - 0 57.66 51.95 64.82 63.37 59.28 68.09 60.76 56.90 65.21 38.49 31.52 49.42 55.85 52.81 59.31
10 53.31 48.11 59.88 60.21 56.86 64.01 53.05 55.76 50.85 44.81 40.58 50.17 53.74 60.24 48.56
20 58.40 55.33 61.86 62.57 60.07 65.32 53.20 54.33 52.13 53.08 50.41 56.29 55.80 55.87 55.79
30 59.50 56.25 63.20 65.79 63.04 68.79 57.19 57.06 57.39 56.07 53.93 58.63 57.97 59.66 56.38
40 58.89 57.23 60.67 67.40 66.97 67.84 59.35 57.90 60.95 60.19 58.85 61.65 57.96 59.63 56.38
50 61.39 60.27 62.57 68.40 68.15 68.65 60.17 56.92 63.84 58.04 56.22 60.00 59.99 62.49 57.71
SC 0 60.56 59.13 62.06 63.87 65.79 62.06 63.56 64.10 63.03 41.48 38.97 44.33 56.57 58.58 54.69
10 48.92 77.35 35.77 48.46 87.50 33.51 43.63 82.78 29.62 15.67 89.29 8.59 42.16 80.13 28.60
20 54.64 83.06 40.71 57.99 87.19 43.44 45.81 82.21 31.75 29.63 86.67 17.87 45.89 91.16 30.66
30 57.55 84.35 43.68 62.61 85.80 49.29 56.14 85.85 41.71 40.32 92.59 25.77 50.25 88.95 35.01
40 59.49 84.67 45.85 62.73 87.34 48.94 58.54 82.05 45.50 38.02 95.83 23.71 52.65 88.17 37.53
50 59.69 86.19 45.65 63.86 88.92 49.82 59.18 82.28 46.21 40.21 91.46 25.77 54.31 89.95 38.90
Opinion Term - 0 52.53 55.87 49.57 57.27 60.43 54.43 53.28 57.82 49.42 29.21 28.16 30.36 39.94 47.49 34.48
10 50.58 48.62 52.72 54.93 50.87 59.70 48.71 49.70 47.76 29.17 26.09 33.08 37.17 43.52 32.44
20 50.39 49.31 51.53 56.71 54.44 59.18 50.90 50.20 51.64 31.01 28.42 34.12 42.90 44.81 41.16
30 50.30 48.96 51.72 59.83 58.21 61.55 55.15 54.85 55.45 33.63 31.13 36.56 43.97 47.00 41.31
40 52.35 50.60 54.25 60.10 58.73 61.55 54.79 54.97 54.64 36.18 34.20 38.42 45.45 49.28 42.18
50 54.40 53.27 55.59 60.31 58.75 61.98 57.14 56.73 57.57 35.96 33.60 38.69 47.89 52.46 44.07
SC 0 52.45 60.82 46.10 56.61 66.18 49.46 55.03 64.83 47.80 29.92 34.41 26.47 40.11 52.98 32.27
10 40.31 75.09 27.55 39.92 75.19 27.17 38.56 78.38 25.57 12.35 68.18 6.79 31.01 65.12 20.35
20 41.98 75.79 29.03 48.92 79.51 35.33 40.99 77.45 27.87 17.19 62.86 9.95 33.59 71.56 21.95
30 46.55 75.76 33.60 53.81 79.16 40.76 49.33 79.30 35.80 26.74 70.19 16.52 39.23 74.69 26.60
40 47.06 74.42 34.41 53.01 78.25 40.08 49.76 76.56 36.86 26.82 75.79 16.29 39.41 72.66 27.03
50 46.41 75.68 33.47 54.17 78.21 41.44 50.77 76.79 37.92 27.59 69.72 17.19 42.01 74.72 29.22
Aspect Category - 0 41.78 45.12 38.91 44.29 46.54 42.25 74.10 77.42 71.08 44.76 46.86 42.85 54.21 61.08 48.75
10 56.62 53.15 60.60 53.13 48.89 58.18 60.97 60.44 61.52 25.64 23.92 27.64 50.45 56.73 45.44
20 60.58 58.76 62.55 58.58 56.02 61.41 64.15 61.38 67.20 32.97 31.64 34.42 53.90 55.40 52.49
30 62.57 61.15 64.07 64.22 62.38 66.18 69.59 67.88 71.40 33.27 31.85 34.83 57.09 59.29 55.06
40 64.31 62.65 66.07 61.71 59.92 63.63 71.64 69.67 73.76 34.74 34.11 35.41 58.36 61.67 55.39
50 64.16 62.96 65.42 65.36 64.40 66.35 71.12 68.66 73.76 34.67 33.70 35.70 58.74 62.20 55.66
SC 0 39.11 45.86 34.10 40.17 46.62 35.29 73.64 82.96 66.20 42.22 52.45 35.33 50.89 62.90 42.74
10 47.45 86.69 32.66 40.89 75.00 28.11 45.60 89.47 30.60 7.18 42.22 3.93 44.28 89.80 29.38
20 50.87 89.53 35.53 49.95 81.46 36.02 50.36 89.74 35.00 16.22 63.38 9.30 45.06 93.19 29.72
30 52.39 85.95 37.68 57.14 84.97 43.05 58.34 90.72 43.00 24.79 69.52 15.08 49.70 90.71 34.22
40 57.14 90.12 41.83 54.11 79.55 41.00 61.38 90.62 46.40 22.45 68.42 13.43 49.64 87.03 34.72
50 54.89 88.25 39.83 57.88 84.31 44.07 63.01 90.94 48.20 24.58 66.36 15.08 51.94 87.70 36.89
Sentiment Polarity - 0 86.90 90.61 83.49 88.05 90.90 85.38 88.68 90.66 86.80 84.41 87.53 81.51 89.57 92.39 86.93
10 84.13 83.12 85.16 84.86 81.11 88.97 85.78 85.59 85.97 85.59 82.76 88.63 81.61 91.41 73.73
20 85.33 82.93 87.90 88.25 86.84 89.72 82.77 79.08 86.85 89.29 88.73 89.85 87.34 91.57 83.49
30 86.98 85.83 88.17 88.12 86.80 89.48 84.05 81.43 86.85 86.41 85.07 87.80 88.52 92.31 85.05
40 87.46 86.57 88.38 88.94 88.33 89.55 87.34 85.57 89.19 88.30 88.03 88.59 88.78 92.99 84.95
50 87.68 87.27 88.10 89.50 88.94 90.07 87.65 85.50 89.93 88.89 88.81 88.98 89.00 93.46 84.95
SC 0 83.29 93.04 75.38 83.35 92.99 75.52 85.45 92.09 79.71 77.82 90.88 68.05 85.23 94.54 77.59
10 56.48 97.08 39.83 53.68 97.29 37.07 53.52 95.60 37.16 17.66 93.02 9.76 56.86 97.70 40.09
20 59.88 97.31 43.25 61.81 97.05 45.34 59.08 96.67 42.54 28.81 100.00 16.83 55.16 97.60 38.44
30 61.95 96.39 45.64 67.19 96.45 51.55 64.72 95.69 48.90 37.25 95.00 23.17 61.61 97.45 45.05
40 65.15 97.61 48.89 65.46 94.77 50.00 68.13 95.18 53.06 36.90 98.94 22.68 64.56 98.08 48.11
50 63.22 96.49 47.01 68.88 96.28 53.62 70.09 96.57 55.01 40.54 97.22 25.61 67.39 97.76 51.42
@@
Table 9: Gemma-2-9B: Performance scores at element-level for the ASQP task. The best score achieved by a method is presented in bold.
@@ Appendix F Performance Scores: Visualization
Figure 4: Influence of the amount of few-shot examples on the performance of Gemma-2-9B and Gemma-2-27B. Visualization includes comparison with performance scores of SOTA supervised methods MVP, Paraphrase and DLO.

### v3

@@ E.2 TASD
Aspect Term - 0 45.84 39.99 53.69 45.26 42.39 48.56 45.49 43.74 47.44 43.44 37.18 52.23 47.20 48.69 45.81
10 50.14 57.58 44.43 50.11 60.08 43.04 50.58 57.47 45.21 59.52 56.31 63.16 52.44 65.20 43.89
20 55.31 61.62 50.22 57.18 69.08 48.79 49.27 58.16 42.84 61.42 61.52 61.37 54.01 64.87 46.27
30 53.39 68.48 43.80 57.33 70.06 48.53 48.49 58.97 41.18 63.34 65.75 61.17 54.84 66.20 46.82
40 53.17 68.42 43.51 54.86 69.46 45.36 49.99 59.04 43.36 65.36 67.50 63.37 60.71 71.67 52.68
50 52.79 68.52 42.95 53.91 70.13 43.79 50.44 58.49 44.36 65.39 68.60 62.47 59.33 72.17 50.39
SC 0 49.34 61.60 41.14 46.17 66.16 35.46 49.63 65.25 40.05 45.54 64.97 35.05 45.19 75.40 32.27
10 40.06 89.17 25.83 44.92 81.20 31.05 50.16 76.96 37.20 59.23 87.84 44.67 47.63 82.95 33.41
20 55.90 85.55 41.51 57.42 85.21 43.30 46.03 76.37 32.94 58.72 88.28 43.99 49.84 86.44 35.01
30 50.00 84.96 35.42 55.39 81.33 41.99 46.43 73.71 33.89 64.22 86.13 51.20 52.02 81.46 38.22
40 50.64 82.50 36.53 53.16 79.74 39.87 51.38 73.25 39.57 66.39 82.23 55.67 62.97 86.75 49.43
50 49.87 79.44 36.35 52.37 80.13 38.89 52.80 76.58 40.28 65.84 82.81 54.64 57.02 80.42 44.16
Aspect Category - 0 41.04 43.15 39.12 51.78 53.50 50.17 57.51 63.16 52.84 30.99 33.11 29.13 40.11 48.67 34.12
10 46.48 49.13 44.10 59.96 63.91 56.47 61.46 68.78 55.56 38.23 39.88 36.74 53.64 64.45 45.94
20 57.59 60.07 55.30 64.90 65.47 64.33 61.61 66.60 57.32 44.36 45.21 43.55 55.32 62.26 49.78
30 61.32 64.58 58.40 69.49 69.43 69.56 64.34 68.59 60.60 45.98 47.03 45.00 60.71 65.74 56.39
40 65.69 68.73 62.92 69.31 69.58 69.10 68.01 72.36 64.16 44.75 45.64 43.93 61.91 66.50 57.93
50 67.06 70.52 63.95 69.71 70.48 68.96 68.88 72.02 66.00 44.53 46.02 43.14 61.05 65.65 57.06
SC 0 34.73 53.64 25.68 45.94 73.02 33.51 51.47 77.20 38.60 26.03 50.30 17.56 34.01 70.90 22.37
10 39.63 67.81 28.00 55.04 84.64 40.78 53.91 82.64 40.00 40.18 69.19 28.31 51.99 93.10 36.06
20 53.80 74.60 42.06 63.78 78.81 53.57 52.21 78.95 39.00 42.51 68.66 30.79 52.86 84.00 38.56
30 58.51 76.89 47.23 68.92 82.67 59.08 56.70 81.04 43.60 46.78 64.26 36.78 59.52 84.62 45.91
40 64.06 81.47 52.77 69.12 80.50 60.57 63.61 83.44 51.40 45.07 59.26 36.36 61.87 80.48 50.25
50 64.51 80.23 53.94 69.44 81.12 60.70 63.70 83.23 51.60 45.99 63.18 36.16 62.01 80.53 50.42
Sentiment Polarity - 0 79.67 82.20 77.31 78.43 80.79 76.22 80.60 85.85 75.99 75.74 80.89 71.22 74.22 79.75 69.43
10 83.94 87.02 81.08 82.67 85.67 79.87 84.04 88.10 80.34 83.15 85.44 80.98 80.60 87.82 74.48
20 87.49 91.04 84.21 87.58 89.26 85.97 83.09 85.83 80.54 85.21 86.25 84.20 88.41 91.18 85.80
30 87.48 90.70 84.49 88.85 90.25 87.49 83.86 86.30 81.56 85.36 86.99 83.80 90.97 93.45 88.63
40 88.09 91.70 84.77 89.29 91.32 87.37 86.99 90.15 84.06 87.25 89.61 85.02 91.14 93.63 88.77
50 89.02 92.58 85.73 89.42 91.17 87.75 86.95 89.33 84.69 85.67 89.39 82.24 91.16 93.99 88.49
SC 0 62.17 91.57 47.06 60.06 91.26 44.76 70.08 95.76 55.26 50.89 94.08 34.88 56.81 96.07 40.33
10 63.48 96.83 47.21 67.56 95.64 52.22 71.74 98.30 56.48 62.02 96.89 45.61 66.67 96.43 50.94
20 75.21 96.14 61.76 80.44 94.83 69.84 70.92 97.44 55.75 63.46 92.52 48.29 73.21 96.53 58.96
30 79.12 96.23 67.18 83.78 95.53 74.60 73.19 95.29 59.41 73.56 93.26 60.73 79.72 96.96 67.69
40 80.43 95.14 69.66 85.81 94.99 78.25 77.76 94.10 66.26 76.92 94.98 64.63 84.57 96.95 75.00
50 82.82 96.11 72.76 86.21 95.03 78.89 79.31 96.17 67.48 71.92 92.02 59.02 84.62 96.67 75.24
@@
Table 9: Gemma-3-4B: Performance scores at element-level for the TASD task. The best score achieved by a method is presented in bold.
@@ Appendix F Performance Scores: Visualization
Figure 4: Influence of the amount of few-shot examples on the performance of Gemma-3-4B and Gemma-3-27B. Visualization includes comparison with performance scores of SOTA supervised methods MVP, Paraphrase and DLO.